OpenStudy (anonymous):

How do i solve for y in this equation: -2x+5y=12

OpenStudy (anonymous):

You have to get the "y" alone one side. so first move the -2x to the other side, so it becomes 5y=12+2x and now you still have that 5 infront of the y, to get rid of that divide both sides by 5 5y/5=(12+2x)/5 you are left with y=(12/5)+(2x/5)